If the voltage V in an electric circuit is held​ constant, the current I is inversely proportional to the resistance R. If the current is 50 amperes when the resistance is 240 ohms, find the current when the resistance is 150




1
Expert's answer
2022-04-08T10:05:59-0400

The voltage is

"V=IR."

If the voltage is constant:

"I_1R_1=I_2R_2,\newline
I_2=I_1\frac{R_1}{R_2}=50\cdot \frac{240}{150}=80\space A."